OpenStudy (anonymous):

make the Graphics r=2cos(3theta)

OpenStudy (tkhunny):

Do you have a plan fior this? Select \(\theta\) from, maybe, 0, (\pi/6\),(\pi/4\),(\pi/3\),(\pi/2\) and get a feel for where it is going. Alternatively, you could convert to Cartesian Cooridnates and get an idea where it's going. Do you know how to do that?